# Estimating the Influence of Sequentially Correlated Literary Properties in Textual Classification: A Data-Centric Hypothesis-Testing Approach

### v2

@@ 50 Potter_2 Kipling_JungleBook2
    Table A.3: All permutations of same-author text pairs used in both experiments. Additionally, the fourth column lists our assumption regarding whether or not each pair of texts varies greatly in genre. When indecisive, we mark that pair with ”?”.
@@ 55 Austen_PP Austen_MansfieldPark ?
    Appendix B Traditional Embeddings Significance Maps
    B.1 Unsupervised Classification Results
@@
    Figure B.1: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
@@
    Figure B.2: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
@@
    Figure B.3: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
    B.2 Supervised Classification Results
@@
    Figure B.4: Significance map for the attempt to apply GI supervised classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
@@
    Figure B.5: Significance map for the attempt to apply GI supervised classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
@@
    Figure B.6: Significance map for the attempt to apply GI supervised classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
@@
    Figure B.7: Significance map for the attempt to apply GI supervised classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
    Appendix C Neural Embeddings Significance Maps
    C.1 Unsupervised Classification Results
    Figure C.1: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using STAR, similarly to Figure 2.
    Figure C.2: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using STAR, similarly to Figure 2.
    C.2 Supervised Classification Results
@@
    Figure C.3: Significance map for the attempt to apply supervised cosine similarity classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using STAR, similarly to Figure 2.
@@
    Figure C.4: Significance map for the attempt to apply supervised cosine similarity classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using STAR, similarly to Figure 2.

### v4

@@ 50 Potter_2 Kipling_JungleBook2
    Table A.4: All permutations of same-author text pairs used in both experiments. Additionally, the fourth column lists our assumption regarding whether or not each pair of texts varies greatly in genre. When indecisive, we mark that pair with “?”.
@@ 55 Austen_PP Austen_MansfieldPark ?
    Appendix B Classification Algorithms with our Hypothesis-Testing Approach
    Algorithm 2 Experiment: Unsupervised ($k$-means)
    1:Initiate MCC score array $v$ of length 100.
    2:Initiate $p$-value array $u$ of length 100.
    3:for $i=1,\ldots,100$ do
    4:    Draw a subsample of $c\cdot\mathrm{min}\{m^{(\ell)}_{1},m^{(\ell)}_{2}\}$ text units ($T^{(\ell,n,f)}_{\mathrm{subsample}}$).
    5:    Apply $2$-means on $T^{(\ell,n,f)}_{\mathrm{subsample}}$ and receive label sequence $L$.
    6:    Compute MCC score between $L$ and $L^{\mathrm{(\mathrm{real})}}$ and store in $v$.
    8:end for
    9:Apply FDR correction to $u$ to receive an FDR-corrected array $u_{\mathrm{FDR}}$.
    10:Chosen MCC score and $p$-value for the combination of $n$ and $\ell$, respectively: $\bar{v}$, $\bar{u}_{\mathrm{FDR}}$.
    Algorithm 3 Experiment: Supervised (GI)
    1:Initiate MCC score array $v$ of length 100.
    2:Initiate $p$-value array $u$ of length 100.
    3:for $i=1,\ldots,100$ do
    4:    Draw a subsample of $c\cdot\mathrm{min}\{m^{(\ell)}_{1},m^{(\ell)}_{2}\}$ text units for training ($T^{(\ell,n,f)}_{\mathrm{train}}$), the rest are for testing ($T^{(\ell,n,f)}_{\mathrm{test}}$).
    5:    Compute two label-wise centroids of $T^{(\ell,n,f)}_{\mathrm{train}}$ according to $L^{\mathrm{(\mathrm{real})}}$.
    6:    Initiate label-association array $w\in\mathbb{R}^{100\times|T^{(\ell,n,f)}_{\mathrm{test}}|}$.
    7:    for $j=1,\ldots,100$ do
    8:        Draw a subsample of $f$, satisfying $f_{\mathrm{subsample}}\in f$ and $|f_{\mathrm{subsample}}|\approx 0.3\cdot|f|$.
    9:        for $k=1,\ldots,|T^{(\ell,n,f)}_{\mathrm{test}}|$ do
    10:           Compute the $2$-norm of the difference between $T^{(\ell,n,f_{\mathrm{subsample}})}_{\mathrm{test},k}$ and each label-wise centroid, respectively.
    11:           $w_{jk}$ = label of the centroid for which the $2$-norm was smaller.
    12:        end for
    13:        Determine the label for every sample in $T^{(\ell,n,f)}_{\mathrm{test}}$ based on the one with which it is most associated in $w$, and receive a label sequence $L$.
    14:    end for
    15:    Compute MCC score between $L$ and $L^{\mathrm{(\mathrm{real})}}$ and store in $v$.
    17:end for
    18:Apply FDR correction $u$ to receive an FDR-corrected array $u_{\mathrm{FDR}}$
    19:Chosen MCC score and $p$-value for the combination of $n$ and $\ell$, respectively: $\bar{v}$, $\bar{u}_{\mathrm{FDR}}$
    Algorithm 4 Experiment: Supervised (Cosine Similarity)
    1:Initiate MCC score array $v$ of length 100.
    2:Initiate $p$-value array $u$ of length 100.
    3:for $i=1,\ldots,100$ do
    4:    Draw a subsample of $c\cdot\mathrm{min}\{m^{(\ell)}_{1},m^{(\ell)}_{2}\}$ text units for training ($T^{(\ell,n,f)}_{\mathrm{train}}$), the rest are for testing ($T^{(\ell,f)}_{\mathrm{test}}$).
    5:    Compute two label-wise centroids of $T^{(\ell,f)}_{\mathrm{train}}$ according to $L^{\mathrm{(\mathrm{real})}}$.
    6:    Initiate label-association array $w\in\mathbb{R}^{|T^{(\ell,f)}_{\mathrm{test}}|}$.
    7:    for $j=1,\ldots,|T^{(\ell,f)}_{\mathrm{test}}|$ do
    8:        $L_{j}=\mathrm{index}[\mathrm{max}(\mathrm{cos}(\mathrm{centroid}_{1},T^{(\ell,f)}_{\mathrm{test},j}),\mathrm{cos}(\mathrm{centroid}_{2},T^{(\ell,f)}_{\mathrm{test},j}))]$.
    9:    end for
    10:    Compute MCC score between $L$ and $L^{\mathrm{(\mathrm{real})}}$ and store in $v$.
    12:end for
    13:Apply FDR correction $u$ to receive an FDR-corrected array $u_{\mathrm{FDR}}$
    14:Chosen MCC score and $p$-value for the combination of $n$ and $\ell$, respectively: $\bar{v}$, $\bar{u}_{\mathrm{FDR}}$
    Appendix C Z-Score Embedding Based on Burrows’ Delta
    To complement traditional tf-idf embeddings, we implement a stylometric embedding inspired by Burrows’ Delta (Burrows, 2002), reformulated for compatibility with clustering algorithms operating in Euclidean space.
    Let $N$ be the number of text units in a given comparison (e.g., blocks of $\ell$ words), and let $f$ denote the number of most frequent $n$-grams selected across the corpus. We construct a document-feature matrix $X\in\mathbb{R}^{N\times f}$, where each entry $x_{ij}$ corresponds to the raw frequency of the $j$th $n$-gram in the $i$th text unit.
    We then standardize $X$ column-wise to obtain the Z-score matrix $Z\in\mathbb{R}^{N\times f}$, whose entries are given by:
    $Z_{ij}=\frac{x_{ij}-\mu_{j}}{\sigma_{j}}$
    where $\mu_{j}=\frac{1}{N}\sum_{i=1}^{N}x_{ij}$ is the mean frequency of feature $j$, and $\sigma_{j}$ is the corresponding standard deviation. In cases where $\sigma_{j}=0$, the column is treated as constant and replaced with zeros in $Z$ to ensure numerical stability.
    This transformation yields a stylometric embedding in which each feature reflects how unusually frequent a given $n$-gram is relative to its corpus-wide baseline. We apply $k$-means clustering directly to the matrix $Z$, using Euclidean distance in the standardized feature space. The resulting cluster assignments are then evaluated against ground truth labels using the Matthews correlation coefficient (MCC), and their statistical significance is assessed via the hypothesis-testing framework described in Section 3.
    Appendix D Traditional Embeddings Significance Maps
    D.1 Unsupervised Classification Results
@@
    Figure D.1: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
@@
    Figure D.2: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
@@
    Figure D.3: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
@@
    Figure D.4: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using word $n$-grams and Z-score embedding with $f$ = 300, similarly to Figure 2.
@@
    Figure D.5: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using character $k$-mers and Z-score embedding with $f$ = 300, similarly to Figure 2.
@@
    Figure D.6: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using word $n$-grams and Z-score embedding with $f$ = 300, similarly to Figure 2.
@@
    Figure D.7: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using character $k$-mers and Z-score embedding with $f$ = 300, similarly to Figure 2.
    D.2 Supervised Classification Results
@@
    Figure D.8: Significance map for the attempt to apply GI supervised classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
@@
    Figure D.9: Significance map for the attempt to apply GI supervised classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
    Figure D.10: Significance map for the attempt to apply GI supervised classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
    Figure D.11: Significance map for the attempt to apply GI supervised classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
    Appendix E Neural Embeddings Significance Maps
    E.1 Unsupervised Classification Results
    Figure E.1: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using STAR, similarly to Figure 2.
    Figure E.2: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using STAR, similarly to Figure 2.
    E.2 Supervised Classification Results
    Figure E.3: Significance map for the attempt to apply supervised cosine similarity classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using STAR, similarly to Figure 2.
    Figure E.4: Significance map for the attempt to apply supervised cosine similarity classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using STAR, similarly to Figure 2.